OpenStudy (anonymous):

What number must be added to the polynomial to complete the square? If necessary, use the slash mark (/)as a fraction bar and leave your answer in the form of an improper fraction. x^2-10x

OpenStudy (anonymous):

10=5=25/4

jimthompson5910 (jim_thompson5910):

how did you go from 5 to 25/4?

jimthompson5910 (jim_thompson5910):

you took half of -10 to get -5 then you're supposed to square -5 to get 25

OpenStudy (anonymous):

oh so -25/4

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

wat?

jimthompson5910 (jim_thompson5910):

half of -10 is ____

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

-5

jimthompson5910 (jim_thompson5910):

then square it to get _____

OpenStudy (anonymous):

-25

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

(-5)^2 = (-5)*(-5) = ??

OpenStudy (anonymous):

+25

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

i wrote the equatoin wrong man its uppose to say x^2-20x

jimthompson5910 (jim_thompson5910):

take half of -20 to get _____

OpenStudy (anonymous):

-10=100? is that the final answer

jimthompson5910 (jim_thompson5910):

yep, 100 is the final answer